OpenStudy (airyana1114):

I'm not sure what I'm solving for? If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = −16t2 + 64t + 4.

OpenStudy (anonymous):

Here h=4 ft from the problem 4ft=-16t^2+64t+4 (devide both sides by 4) 1ft=-4t^2+16t+1 (can not be factored, solve by quadratic equation)

OpenStudy (phi):

how long will it take the baseball to hit the ground? The ground is assumed to be at height 0. they are asking for what time "t" makes this equation true: \[ −16t^2 + 64t + 4 =0\] you can simplify that a little by dividing both sides by by -4 (that means divide all terms by -4). You get \[ 4t^2 -16t-1=0 \] See https://www.khanacademy.org/math/algebra/quadratics/quadratic-formula/v/quadratic-formula-1 for an example of how to use the quadratic formula to solve for t you will get two numbers, one negative (which is "back in time") and one positive number. The answer will be the positive number